OpenStudy (anonymous):

what is the boiling point of a 0.75m water solution? water has a boiling point elevation of 0.52 degrees celcious/m.

OpenStudy (anonymous):

you must find the increasing of boiling point \[\Delta Tb = molal \times Kb\] = 0.75 x 0.52 = 0.39 Then the boiling point = 100 + 0.39 = 100.39 degree Celcius